Raymond Mining Corporation has 10.1 million shares of common stock outstanding, 450,000 shares of 5% $100 par value preferred stock outstanding, and 175,000 7.50% semiannual bonds outstanding, par value $1,000 each. The common stock currently sells for $49 per share and has a beta of 1.55, the preferred stock currently sells for $98 per share, and the bonds have 15 years to maturity and sell for 116% of par. The market risk premium is 8.9%, T-bills are yielding 4%, and Raymond Mining’s tax is 38%. The YTM before taxes is 5.880%.

What is the WACC? (Enter your answers in whole dollars. Omit % sign in your response.)
Answer #1

Value of Common Share = 10.1 mn shares * 49 = 494.9 mn

Value of Preferred Share = 450,000 * 98 = 44.1 mn

Value of Debt = 750,000 * 1000 * 116% = 870 mn

Total Capital = 494.9 + 44.1 +870 = 1409 mn

Cost of Equity = Rf + beta * (Rm - Rf)

= 4% * 1.55 * 8.9%

= 17.80%

Cost of Preferred Equity = Dividend / Cost of Equity  

= 5 / 98

= 5.10%

Cost of Debt = 5.88%

WACC = (Cost of Equity * Weight of Equity) + (Cost of Debt after tax * Weight of Debt) + (Cost of Preferred Stock * Cost of Preferred Stock)

= 17.80% * 494.9 / 1409 * 5.88 * (1-0.38) * 870 / 1409 + 5.10% * 44.1 / 1409

= 8.66%